# https://rlwg-grndctrl.github.io/

### v1

@@ 7 Conclusions
Our work demonstrates that RLWG enables pretrained video world models to be effectively grounded in physical structure through self-supervised post-training, without requiring human labels or external simulators. Instantiated as GrndCtrl, the key insight is that verifiable geometric and perceptual rewards, when optimized via relative policy methods, systematically improve spatial coherence while preserving the visual quality that makes these models powerful priors. The most significant improvements emerge under counterfactual scenarios. GrndCtrl achieves substantial improvements under counterfactual rollouts, with up to $64\%$ reduction in translation error, demonstrating that reward-based alignment addresses a fundamental limitation of current world models: their tendency to prioritize visual plausibility over structural consistency. By explicitly optimizing for both, RLWG enables models that are not only visually coherent but also geometrically consistent, opening new possibilities for reliable long-horizon planning and control in real-world environments.
@@ Acknowledgements
We thank NVIDIA and the authors of Cosmos-Predict2 [nvidia2025cosmosworldfoundationmodel] for publicly releasing their code and checkpoints, upon which we build our post-training pipeline. We thank Nikhil Keetha and collaborators for MapAnything [keetha2025mapanythinguniversalfeedforwardmetric], and the authors of VideoAlign [liu2025improving], for making their code and checkpoints publicly available, enabling our verifiable self-supervised rewards and evaluation metrics. We also thank Amir Bar and collaborators for releasing Navigation World Models [bar2025navigationworldmodels] code and checkpoints and for kindly responding to our inquiry email. Finally, we thank Cherie Ho for early-stage discussions and Jay Karhade for discussions on geometric consistency.
[…]
@@ Appendix A Qualitative Results
We include additional counterfactual generation results of GrndCtrl against baseline and ground truths in SCAND [karnan2022sociallycompliantnavigationdataset] (Figure 4) and CityWalk [liu2025citywalker] (Figure 5), in addition to CODa [zhang2023towards] (Figure 3). The counterfactual rollouts of GrndCtrl and baseline are expected to mirror the trajectory shown in ground truths in each scene. We observe GrndCtrl improves trajectory-following and mitigates scene drift in counterfactual rollouts in all three datasets.
@@
Appendix B Comparisons with NWM [bar2025navigationworldmodels]
While RLWG is a post-training framework agnostic to the baseline world model, and in GrndCtrl we obtain the baseline with supervised fine-tuning, we acknowledge other potential candidates that perform similar tasks as an experimental baseline. This is exemplified by NWM [bar2025navigationworldmodels], which operates a similar world models scenario in navigation tasks. However, significant differences in experimental settings between GrndCtrl and NWM results in infeasible quantitative comparisons:
  • NWM is jointly-trained on multiple datasets, including RECON [shah2021rapid] and HuRoN [hirose2023sacson], which are collected using fish-eye cameras. This results in frequent twisted artifacts in rollouts conditioned on rectilinear images such as SCAND [karnan2022sociallycompliantnavigationdataset]. This biases NWM to perform poorly with visual odometry metrics based on a feed-forward 3D evaluator such as MapAnything [keetha2025mapanythinguniversalfeedforwardmetric].
@@ D.4 Evaluators and Rewards
MapAnything [keetha2025mapanythinguniversalfeedforwardmetric] serves as our frozen 3D evaluator $E$, providing relative pose estimates $(\Delta R_{t},\Delta\mathbf{t}_{t})$ and per-frame depth maps $D_{t}$ for each rollout $\hat{x}_{1:T}$. VideoAlign [liu2025improving] serves as our frozen video evaluator $V$, providing sequence-level visual quality, motion quality, and text alignment scores. We use only visual and motion quality scores (with equal weighting $\alpha=\beta=0.5$) for the video quality reward $r_{\text{v}}$, ignoring text alignment since our rollouts are not text-conditioned. When GPU memory is constrained, model parameters of the evaluators are temporarily offloaded to CPU, and re-loaded only when used.

### v2

@@ 7 Conclusions
Our work demonstrates that RLWG enables pretrained video world models to be effectively grounded in physical structure through self-supervised post-training, without requiring human labels or external simulators. Instantiated as GrndCtrl, the key insight is that verifiable geometric and perceptual rewards, when optimized via relative policy methods, systematically improve spatial coherence while preserving the visual quality that makes these models powerful priors. The most significant improvements emerge under counterfactual scenarios. GrndCtrl achieves substantial improvements under counterfactual rollouts, with up to $64\%$ reduction in translation error, demonstrating that reward-based alignment addresses a fundamental limitation of current world models: their tendency to prioritize visual plausibility over structural consistency. By explicitly optimizing for both, RLWG enables models that are not only visually coherent but also geometrically consistent, opening new possibilities for reliable planning and control in real-world environments.
@@ Acknowledgements
We thank NVIDIA and the authors of Cosmos-Predict2 [17] for publicly releasing their code and checkpoints, upon which we build our post-training pipeline. We thank Nikhil Keetha and collaborators for MapAnything [13], and the authors of VideoAlign [15], for making their code and checkpoints publicly available, enabling our verifiable self-supervised rewards and evaluation metrics. We also thank Amir Bar and collaborators for releasing Navigation World Models [4] code and checkpoints and for kindly responding to our inquiry email. Finally, we thank Cherie Ho for early-stage discussions and Jay Karhade for discussions on geometric consistency.
@@ Appendix A Qualitative Results
We include additional counterfactual generation results of GrndCtrl against baseline and ground truths in SCAND [10] (Figure 4) and CityWalk [16] (Figure 5), in addition to CODa [28] (Figure 3). The counterfactual rollouts of GrndCtrl and baseline are expected to mirror the trajectory shown in ground truths in each scene. We observe GrndCtrl improves trajectory-following and mitigates scene drift in counterfactual rollouts in all three datasets.
@@
Appendix B Comparisons with NWM [4]
While RLWG is a post-training framework agnostic to the baseline world model, and in GrndCtrl we obtain the baseline with supervised fine-tuning, we acknowledge other potential candidates that perform similar tasks as an experimental baseline. This is exemplified by NWM [4], which operates a similar world models scenario in navigation tasks. However, significant differences in experimental settings between GrndCtrl and NWM results in infeasible quantitative comparisons:
  • NWM is jointly-trained on multiple datasets, including RECON [20] and HuRoN [8], which are collected using fish-eye cameras. This results in frequent twisted artifacts in rollouts conditioned on rectilinear images such as SCAND [10]. This biases NWM to perform poorly with visual odometry metrics based on a feed-forward 3D evaluator such as MapAnything [13].
@@ D.4 Evaluators and Rewards
MapAnything [13] serves as our frozen 3D evaluator $E$, providing relative pose estimates $(\Delta R_{t},\Delta\mathbf{t}_{t})$ and per-frame depth maps $D_{t}$ for each rollout $\hat{x}_{1:T}$. VideoAlign [15] serves as our frozen video evaluator $V$, providing sequence-level visual quality, motion quality, and text alignment scores. We use only visual and motion quality scores (with equal weighting $\alpha=\beta=0.5$) for the video quality reward $r_{\text{v}}$, ignoring text alignment since our rollouts are not text-conditioned. When GPU memory is constrained, model parameters of the evaluators are temporarily offloaded to CPU, and re-loaded only when used.